OpenStudy (anonymous):

Evaluate the indefinite integral of: 1 ---- 1 + (t)^1/2 dt I would like a hint. I know substitution is involved.

OpenStudy (zarkon):

\[\frac{1}{1+\sqrt{t}}=\frac{1}{\sqrt{t}}-\frac{1}{\sqrt{t}(\sqrt{t}+1)}\]

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

I was able to get u = (t)^1/2, but then I could not find an expression for 1/2(t)^1/2

OpenStudy (zarkon):

you could have also let \[u=\sqrt{t}\] and worked it out from there

OpenStudy (zarkon):

\[du=\frac{1}{2\sqrt{t}}dt\] \[2\sqrt{t}du=dt\] \[2udu=dt\] \[\int\frac{1}{1+\sqrt{t}}dt=\int\frac{2u}{1+u}du\]

OpenStudy (zarkon):

\[=2\int\frac{u}{1+u}du=2\int\frac{1+u-1}{1+u}du\] \[=2\int\left[\frac{1+u}{1+u}+\frac{-1}{1+u}\right]du=2\int\left[1+\frac{-1}{1+u}\right]du\] \[=2(u-\ln(|1+u|)+c=2\sqrt{t}-2\ln(\sqrt{t}+1)+c \]